CH_{4}/CO_{2} REFORMING OVER La_{2}NiO_{4} AND 10 PERCENT NiO/CeO_{2}-La_{2}O_{3} CATALYSTS UNDER THE CONDITION OF SUPERSONIC JET EXPANSION VIA CAVITY RING DOWN SPECTROSCOPIC ANAYL-SIS

B.S. LIU, Department of Chemistry, Tianjin University Tianjin 300072, P.R. China and Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong; J. W-H. LEUNG, LING LI AND \underline{A. S-C. CHEUNG}, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong; C.T. AU, Department of Chemistry and Center for Surface Analysis and Research, Hong Kong Baptist University, Kowloon Tong, Hong Kong.

The catalytic reaction of CO_{2} reforming of CH_{4} produces synthesis gas (CO and H_{2}), which is useful as a feedstock for many industrial processes. In the past few years, many transition metal-based catalysts have been investigated but the major problems reported were catalytic deactivation due to carbon deposition and low yield. We studied the CH_{4}/CO_{2} reforming over La_{2}NiO_{4} and 10%NiO/CeO_{2}-La_{2}O_{3} catalysts under the condition of supersonic jet expansion via direct monitoring of the change in reactants (CH_{4} and CO_{2}), product (CO) and side-product (H_{2}O) using the sensitive technique of cavity ring-down spectroscopy. Vibration-rotational absorption lines of CH_{4}, H_{2}O, CO_{2} and CO molecules were recorded in the near infrared spectral region. We found that La_{2}NiO_{4} is superior to 10%NiO/CeO_{2}-La_{2}O_{3} in performance. We have also investigated the associated reverse water-gas shift (RWGS) reaction, which affects significantly the H_{2}/CO product ratio, over the catalysts during CH_{4}/CO_{2} reforming. Our results indicated that the RWGS reaction promoted the conversion of CO_{2} and decreased the partial pressure of hydrogen. By proper adjustment of the pressure of the reaction system, it is possible to suppress the occurrence of RWGS reaction and increase the selectivity of CH_{4}/CO_{2} reforming.